OpenStudy (anonymous):

6.) the following function represents the profit P(n) in dollars that a concert promoter makes by selling tickets for n dollars each: P(n)=-250n^2+3,250n-9,000 Part A: What are the zeroes of the above function and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n ). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n )?

OpenStudy (valpey):

You can see that as n gets very large, profit becomes negative. Profit is also negative when n is zero. The nature of this second degree (n^2) polynomial is a parabola. It has only one maximum when the derivative w.r.t. n is equal to zero. This occurs when 0 = -500n+3250

OpenStudy (anonymous):

@Compassionate @tHe_FiZiCx99 @Zale101

hero (hero):

You are given the function P(n) = -250n^2+3,250n-9,000.

hero (hero):

In Part A you are asked to find the zeroes of the function, right?

OpenStudy (anonymous):

oh I got part a already

OpenStudy (anonymous):

I factored 0 = -250n^2+3,250n-9,000 and got -250(n-4) (n-9) which is the zero’s or solution of the function which represents the x intercepts.

hero (hero):

Okay, Hang on, let me check that

OpenStudy (sarahezzmck):

is this the one you asked for my help on?

hero (hero):

Okay, that looks good. In Part B, you are asked to find the maximum profit by completing the square, right?

OpenStudy (anonymous):

yes

hero (hero):

Okay, so to complete the square, we have to first put P(n) in the form Ax^2 + Bx = C right?

OpenStudy (anonymous):

yeah

hero (hero):

The first step is to Set P(n) = 0 so 0 = -250n^2+3,250n-9,000

hero (hero):

Then add 9000 to both sides to get 9000 = -250n^2 + 3250n which is the same as -250n^2 + 3250n = 9000 Notice that has the same form as Ax^2 + Bx = C

hero (hero):

Actually, if we divide both sides by -1 we get 250n^2 - 3250n = -9000 Now we have it in the form Ax^2 + Bx = C

hero (hero):

Now we need to divide both sides by 250. After doing so we get n^2 - 13n = -36

hero (hero):

At this point we have it in a form where we can complete the square. Do you agree?

OpenStudy (anonymous):

yes

hero (hero):

Do you remember what needs to be done in order to complete the square? What should we add to both sides?

OpenStudy (anonymous):

you add a needed value to each side to balance the equation out @Hero

hero (hero):

Once you have it in the form x^2 + bx = c, you complete the square by adding \(\left(\frac{b}{2}\right)^2\) to both sides.

hero (hero):

In this case, b = -13 so we will add \(\left(\frac{-13}{2}\right)^2\) to both sides of \(n^2 - 13n = -36\).

hero (hero):

Upon doing so we will have: \(n^2 - 13n + \left(\frac{-13}{2}\right)^2 = \left(\frac{-13}{2}\right)^2 - 36\)

hero (hero):

Do you agree?

OpenStudy (anonymous):

yeah

hero (hero):

Do you know what that simplifes to?

OpenStudy (anonymous):

simplify (-13/2)^2-36 to 25/4?

OpenStudy (anonymous):

right @Hero

hero (hero):

Yes, you correctly simplified the right side. What about the left side?

hero (hero):

We need to write the left side as a binomial square.

OpenStudy (anonymous):

n^2-13n- (13/2)?

hero (hero):

Do you remember how to expand\(\left(\frac{-13}{2}\right)^2\)?

OpenStudy (anonymous):

nope

hero (hero):

Hint: \(\left(\dfrac{a}{b}\right)^2 = \dfrac{a^2}{b^2}\)

hero (hero):

When you expand it, you get 169/4. That was correct.

OpenStudy (anonymous):

oh okay wasn't for sure

hero (hero):

However, I remembered, that you don't need to expand it on the left side because basically \(n^2 - 13n + \left(\frac{-13}{2}\right)^2\) = \(\left(n - \frac{13}{2}\right)^2\)

hero (hero):

So what you end up with is \(\left(n - \frac{13}{2}\right)^2 = \frac{25}{4}\)

hero (hero):

We have to keep solving for n, so what do you think we should to next?

OpenStudy (anonymous):

well I know we have to iscolate the variable so im thinking we subtract from the side with the variable

OpenStudy (anonymous):

@Hero

hero (hero):

Actually, first we have to take the square root of both sides: \(\sqrt{\left(n - \frac{13}{2}\right)^2} = \sqrt{\frac{25}{4}}\)

hero (hero):

What do you know about the relationship between a square and square root?

OpenStudy (anonymous):

not much, ive just heard of them that's it

OpenStudy (anonymous):

why @Hero

OpenStudy (anonymous):

n= 4 and 9 @Hero

OpenStudy (anonymous):

is that right

hero (hero):

You solved for n the rest of the way? Can you show the steps you took to find n?